OpenStudy (anonymous):

The sum of three consecutive integers is 312. What are the integers

OpenStudy (anonymous):

Consecutive just means one after the other (example 1, 2, 3 are consecutive numbers) If you use x to represent the smallest integer, then you can use x + 1 for the number just after it, and x + 2 for the third number. So: x + (x+1) + (x+2) = 312 x + x + 1 + x + 2 = 312 3x = 309 x = 103 Now that you've found the value of the smallest integer you can get the other two without having to do much work. x = 103 x + 1 = 103 + 1 = 104 x + 2 = 103 + 2 = 105 103 + 104 + 105 = 312.